40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 13637     Entry time: Fri Feb 16 15:57:58 2018     In reply to: 13423
Author: Steve 
Type: Update 
Category: VAC 
Subject: TP3 drypump replaced 

The forline pressure of TP3 was 399 mTorr

It was replaced this morning at  TP3 controller 134,638hrs with the "failed TP2 station" drypump. The foreline pressure now at TP3 is 100 mTorr at 6 hrs of operation.[ at day 3  63 mT ]

IFO pressure at CC Hornet 7.9e - 6 Torr

Valve configuration: vacuum normal as TP3 is the forepump of the Maglev & the annuloses are not pumped

Quote:

PSL shutter closed at 6e-6 Torr-it    

   The foreline pressure of the drypump is 850 mTorr at 8,446 hrs of seal life

V1 will be closed for ~20 minutes for drypump replacement..........

9:30am dry pump replaced, PSL shutter opened at 7.7E-6 Torr-it

  Valve configuration: vacuum normal as  TP3 is the forepump of the Maglev  & annuloses are not pumped.

Quote:

TP3 drypump replaced at 655 mTorr, no load, tp3 0.3A 

This seal lasted only for 33 days at  123,840 hrs

The replacement is performing well: TP3 foreline pressure is 55 mTorr, no load, tp3 0.15A at 15 min  [ 13.1 mTorr at d5 ]

 

Valve configuration: Vacuum Normal, ITcc 8.5E-6 Torr

Quote:

Dry pump of TP3 replaced after 9.5 months of operation.[ 45 mTorr d3 ]

The annulosses are pumped.

Valve configuration: vac normal, IFO pressure 4.5E-5 Torr [1.6E-5 Torr d3 ] on new ITcc gauge, RGA is not installed yet.

Note how fast the pressure is dropping when the vent is short.

Quote:

IFO pressure 1.7E-4 Torr on new not logged cold cathode gauge. P1 <7E-4 Torr

Valve configuration: vac.normal with anunulossess closed off.

TP3 was turned off with a failing drypump. It will be replaced tomorrow.

All time stamps are blank on the MEDM screens.

 

 

 

ELOG V3.1.3-